OpenStudy (anonymous):

Square ABCD has side length equal to 12. E is the midpoint of AD and F is the midpoint of DC. What is the area of quadrilateral BEFC?

ganeshie8 (ganeshie8):

start by drawing the square and labeling things

OpenStudy (anonymous):

but wats the ans

ganeshie8 (ganeshie8):

you solve it and see what u get :)

OpenStudy (anonymous):

but m not getting the correct ans

OpenStudy (nirmalnema):

 area of square = a^2 =12*12 =144sq units now area of quadrilateral= area of sq. - area of triangle..AEB and EDF 144-(36+18)=90 sq units